OpenStudy (anonymous):

factor a^2-9a show steps

OpenStudy (anonymous):

Can you identify a common factor in the two terms?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Great. Are you able to factor out the a?

OpenStudy (anonymous):

\[=a^2-9a\]\[=a(a-9)\]\[=?\] solve it @hectorl

OpenStudy (anonymous):

@hector1